# Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

### v3

@@ Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety. Existing detection methods often depend on predefined patterns, limiting their adaptability across diverse LLM architectures. We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by leveraging entropy to quantify prediction uncertainty and a local search strategy for exploring the token space. Experiments across multiple LLM architectures show that GlitchMiner outperforms existing methods in both detection accuracy and adaptability, achieving over 10% average efficiency improvement. GlitchMiner enhances vulnerability assessment in LLMs, contributing to more robust and reliable applications. Code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 . However, the increasing reliance on these models in critical domains has brought to light a subtle yet significant vulnerability: glitch tokens geiping2024coercing . These anomalous tokens can severely disrupt an LLM’s ability to generate coherent and reliable outputs, potentially leading to repetitive errors, nonsensical responses, or even harmful content. The implications of such disruptions are particularly concerning in high-stakes applications where model reliability is paramount.
@@
Recent research efforts have made strides in understanding and detecting glitch tokens. Approaches like Magikarp Magikarp and GlitchHunter GlitchHunter have provided valuable insights, focusing on abnormal embedding patterns and clustering properties of these tokens. However, these methods are inherently limited by their reliance on manually observed patterns and predefined distribution assumptions. This dependence not only constrains their adaptability across diverse LLM architectures but also potentially overlooks novel forms of glitch tokens that do not conform to expected patterns.
@@
In this paper, we introduce GlitchMiner, a novel framework that fundamentally reimagines the approach to glitch token detection. Moving beyond pattern-based heuristics, GlitchMiner employs gradient-based discrete optimization to directly analyze an LLM’s prediction behavior. At the core of our method is the use of entropy as an exploratory metric, guiding our search through the vast token space to identify regions of unexpected model behavior.
In summary, the key innovations of GlitchMiner include:
@@
  • Entropy-Based Optimization: We introduce entropy as a novel optimization objective for glitch token detection. This approach allows us to systematically identify tokens that induce high uncertainty in the model’s predictions, moving beyond predefined patterns or embedding anomalies. By focusing on the model’s prediction behavior, we can detect a broader spectrum of glitch tokens, including those that might be missed by traditional embedding-based methods.
@@
  • Gradient-Guided Local Search: Our method combines first-order Taylor approximation with a local search strategy, efficiently navigating the vast token space. This approach significantly improves both the efficiency and accuracy of glitch token detection, allowing us to explore the token space more thoroughly while maintaining computational feasibility.
@@
  • Adaptive Cross-Architecture Detection: GlitchMiner’s framework is designed to adapt to the unique characteristics of different LLM architectures without relying on model-specific assumptions. This flexibility enables effective glitch token detection across a wide range of models, addressing a key limitation of previous methods that often relied on architecture-specific heuristics.
@@
We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families. The results show that GlitchMiner surpasses existing methods in 90% of cases when it comes to glitch token detection efficiency. Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1. These results represent improvements of 10.5% and 10.7% respectively over the previous leading methods. In the context of glitch token detection, these improvements are significant as they translate to hundreds of additional potentially harmful tokens being identified. Such enhancements not only improve detection performance but also optimize query efficiency, which is crucial for enhancing the reliability and robustness of LLMs in critical applications where even a single undetected glitch token could lead to severe consequences.
@@
The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work in glitch token detection and gradient-based optimization techniques. Section 3 details the design principles and implementation of GlitchMiner. Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures. Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research in this critical area of LLM security and reliability.
@@ 2.1 What is a Glitch Token?
  • They often have abnormal embedding patterns, such as unusually large L2 norms or atypical positions in the embedding space Magikarp .
@@ 3.3 Token Selection
Motivation for Entropy-Based Approach. Previous methods for glitch token detection often rely on manually observed patterns or predefined distribution assumptions. However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens. To address these limitations, we propose an entropy-based approach. Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs. By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
@@ 4.3 Ablation Study
To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, and batch size $B$.

### v4

# GlitchMiner: Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

@@ Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety. Existing detection methods rely on predefined patterns, limiting their adaptability across diverse LLM architectures. We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by introducing entropy as a measure of prediction uncertainty and employing a local search strategy to explore the token space. Experiments across multiple LLM architectures demonstrate that GlitchMiner outperforms existing methods in detection accuracy and adaptability, achieving over 10% average efficiency improvement. This method enhances vulnerability assessment in LLMs, contributing to the development of more robust and reliable applications. Code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 . However, with the growing reliance on these models in critical domains, a subtle yet significant vulnerability has emerged: glitch tokens geiping2024coercing . These anomalous tokens can severely disrupt an LLM’s output, leading to repetitive errors, nonsensical responses, or even harmful content. Such disruptions are particularly concerning in high-stakes applications where model reliability is crucial.
@@
Recent research has made progress in detecting glitch tokens. Methods such as Magikarp Magikarp and GlitchHunter GlitchHunter focus on abnormal embedding patterns and clustering properties. However, these approaches are limited by their dependence on predefined patterns, restricting their adaptability across different LLM architectures and potentially missing novel forms of glitch tokens that deviate from known patterns.
@@
In this paper, we introduce GlitchMiner, a novel framework for glitch token detection that overcomes the limitations of pattern-based heuristics by employing gradient-based discrete optimization to directly analyze a model’s prediction behavior. Central to our method is the use of entropy as an exploratory metric, guiding the search through the token space to identify regions indicative of unexpected model behavior.
The key innovations of GlitchMiner include:
@@
  • Entropy-Based Optimization: By introducing entropy as a novel objective for glitch token detection, we systematically identify tokens that induce high uncertainty in model predictions. This approach enables detection of a broader range of glitch tokens, extending beyond traditional pattern- or embedding-based methods.
@@
  • Gradient-Guided Local Search: Combining first-order Taylor approximation with a local search strategy, GlitchMiner efficiently explores the vast token space. This significantly enhances both the efficiency and accuracy of glitch token detection, allowing thorough exploration while maintaining computational feasibility.
@@
  • Adaptive Cross-Architecture Detection: GlitchMiner is designed to adapt to the unique characteristics of various LLM architectures without relying on model-specific assumptions, enabling effective glitch token detection across a wide range of models.
@@
We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families. The results show that GlitchMiner surpasses existing methods in detection efficiency in 90% of cases. Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1, with respective improvements of 10.5% and 10.7% over state-of-the-art methods. These gains enhance the detection of potentially harmful tokens, improving both accuracy and query efficiency. Such advancements are essential for LLM reliability, especially in applications where undetected glitch tokens could have serious consequences.
@@
The remainder of this paper is structured as follows: Section 2 provides a review of related work in glitch token detection and gradient-based optimization techniques. Section 3 details the design principles and implementation of GlitchMiner. Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures. Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research.
@@ 2.1 What is a Glitch Token?
  • They often have abnormal embedding patterns, such as unusually small $\ell_{2}$ norms or atypical positions in the embedding space Magikarp ; solidgoldmagikarp .
@@ 3.3 Token Selection
Motivation. Previous methods for glitch token detection often rely on manually observed embedding patterns. However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens. To address these limitations, we propose an entropy-based approach. Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs. By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
@@ 4.3 Ablation Study
To evaluate the contributions of key components in GlitchMiner, we conducted ablation studies focusing on the local search strategy, neighborhood size $K$, batch size $B$, and initialization token.